A chemical plant produces an ammonia-rich waste ga
Answer #1

Given : Volumetric flow rate of ammonia = 105 m3/hr = 7/240 m3/s

Mole density of waste gas, dw = 0.0407 mol/L = 40.7 mol/m3

Initial mole fraction of ammonia in incoming waste gas = 0.120

Thus, mole density of ammonia in incoming waste gas = 0.120*40.7 = 4.884 mol/m3

Thus, rate at which ammonia comes in from the incoming stream = mole density of NH3*Volumetric flow rate = 4.884*(7/240) = 0.14245 moles NH3 per second

Since 90% ammonia gets removed, so rate of removal of ammonia = 0.9*rate at which ammonia comes in = 0.9*0.14245 = 0.128 moles NH3 per second = 460.8 mol/h

(2) Ammonia holding capacity of water = 0.35 mol NH3 per mole of water

So, to hold 0.128 moles ammonia, moles of water required = 1/0.35*0.128 = 0.365 moles water

Molecular mass of water = 18 g

So, mass of water required = 18*0.365 = 6.57 g

Thus, volumetric flow rate of water required = mass required/density of water = 6.57/0.99 = 6.63 ml per second = 6.63*10-6*3600 = 0.023868 m3/hr = 23.868 L/h

(3) New mole fraction of ammonia = moles of ammonia in the outgoing stream/total moles in outgoing stream

Moles of ammonia in outgoing stream = 10% of moles of ammonia in incoming stream (because 90% gets removed ) = 0.10*initial mole density = 0.1*4.884 = 0.4884 moles

Total moles in outgoing stream = total moles in initial stream-moles of ammonia removed = 40.7-(90% of 4.884) = 36.3

Thus, new mole fraction of ammonina = 0.4884/36.3 = 0.01345